# Offline Dedicated Quantum Attacks on Block Ciphers Based on Two Parallel Permutation-Based Pseudorandom Functions

### v2

@@ Abstract
Quantum cryptanalysis is essential for evaluating the security of cryptographic systems against the threat of quantum computing. Recently, Shi et al. introduced the dedicated quantum attack on XOR-type function that greatly reduces the required resources (including circuit depth, width, and the number of gates) compared to the parallel Grover-meets-Simon algorithm. Here, our contribution is in two aspects. On the one hand, we discover new cryptographic structures amenable to this attack: PolyMAC and block ciphers based on two parallel permutation-based pseudorandom functions (TPP-PRFs), including $\mathsf{XopEM}$, $\mathsf{SoEM22}$, $\mathsf{SUMPIP}$, and $\mathsf{DS\text{-}SoEM}$, partially answering Shi et al.’s open question. On the other hand, for block ciphers based on TPP-PRFs, we break the obstacle that this attack rely on online query by constructing decoupled XOR-type function, then propose an offline quantum attack on them that retains the tunable truncation parameter, $t$, a positive integer. Compared to previous results, our offline attack exhibits significantly reduced query complexity. Specifically, we reduce the number of queries to the encryption oracle from $\tilde{O}(2^{(n+t)/2})$ to $O(2^{t})$ with the same time complexity in the quantum query model, and enable its implementation in the classical query model, optimizing both the classical query complexity and time complexity from $\tilde{O}(2^{2n/3})$ to $\tilde{O}(2^{(2n-t)/3})$.
[…]
1 Introduction
Quantum computing, rooted in the principles of quantum mechanics, has potential speed-up advantages for solving specific classes of problems, such as combinatorial optimization [33, 35], image classification [32, 22], quantum architecture search  [28, 10], and so on [27, 19]. The rapid progress in quantum computing poses unprecedented threats to cryptographic security. Notably, Shor’s algorithm [25] and Grover’s algorithm [11] fundamentally threaten public-key and symmetric cryptography, respectively.
@@ 1 Introduction
Simon’s algorithm [26] provides a crucial tool for quantum cryptanalysis by achieving exponential speedup over classical methods in finding the period of two-to-one periodic functions. This capability directly threatens the security of block ciphers (e.g., $3$-round Feistel [16], Even-Mansour cipher [17, 24, 34], etc.). In 2017, Grover-meets-Simon algorithm, which is presented by Leander and May [18], demonstrated that key whitening does not enhance security in the quantum-chosen-plaintext attack (q-CPA) setting on the FX construction. Subsequently, Grover-meets-Simon algorithm has played a pivotal role in researching the quantum security of symmetric cryptography, including $\mathsf{Feistel}$ [9, 14], $\mathsf{Lai\text{-}Massey}$ [20], $\mathsf{MISTY}$ structure [6], $\mathsf{SUM\text{-}ECBC}$ [13], and several Beyond-Birthday-Bound MACs [29].
@@
The above attack relies on superposition queries to the encryption oracle (the Q2 model), which present significant implementation challenges. This has spurred growing interest in the attack that utilizes classical query and offline quantum computation (the Q1 model) in recent years. In 2019, Bonnetain et al. [4] proposed the offline Simon’s algorithm, dividing it into two distinct phases: the online phase for accessing the encryption oracle to prepare the superposition state and the offline computation phase for running the Grover iteration on this state to verify whether the function exhibits periodicity. They achieved key-recovery attacks on the FX and Even-Mansour constructions in the Q1 model, significantly optimizing attack complexities. In 2021, Bonnetain [3] explored the quantum implementation of the offline Simon’s algorithm. Further, Li et al. [19] proposed a quantum key recovery attack on $\mathsf{SoEM21}$ and $\mathsf{SoEMs1}$ in the Q1 model. Then, Sun et al. [31] systematically studied the quantum security of permutation-based pseudorandom functions, where the adversaries are restricted to classical queries and offline quantum computations.
@@
As the National Institute of Standards and Technology points out, due to the limitations of quantum computer implementation technology, reducing the circuit depth of quantum attacks is essential for their practical implementation. Shi et al. [23] proposed a dedicated quantum attack on XOR-type function, which is written in the following form: $f(\alpha_{k},x)=g_{1}(x)\oplus g_{2}(\alpha_{k}\oplus x)$. In contrast to the parallelization of Grover-meets-Simon algorithm, their attack saved on circuit costs (e.g., the depth, width, and gate count), while requiring a quantum query complexity of $2^{(\kappa+t)/2}$ to the encryption oracle. Moreover, they analyzed that the attack has enabled improved cryptanalysis of several Beyond-Birthday-Bound (BBB) MACs, such as $\mathsf{SUM\text{-}ECBC}$, its variants $2\mathsf{K}$-$\mathsf{SUM\text{-}ECBC}$, $2\mathsf{K}$-$\mathsf{ECBC\_Plus}$, and $\mathsf{GCM\text{-}SIV2}$.
@@
Motivations. Our work is motivated by two critical open problems arising from the preceding discussion.
@@
First, the dedicated quantum attack proposed by Shi et al. [23] has only been applied to the BBB MACs mentioned above. A open question posed by the authors is to identify other cryptographic constructions that are vulnerable to this attack.
@@
Second, this attack on these BBB MACs relies on superposition queries to the encryption oracle. This naturally leads to a more challenging question. Can the quantum attack on XOR-type function be performed in the Q1 model and other instances of block ciphers?
@@
Our Contributions. Addressing these challenges of applicability and practicality, our work achieves the following breakthroughs:
@@
    we construct novel XOR-type function based on PolyMAC and two parallel permutation-based pseudorandom functions (TPP-PRFs). These constructions successfully provide practical instances of XOR-type function, thereby answering the open question posed by Shi et al. [23].
@@
    we propose the offline dedicated quantum attack on block ciphers based on TPP-PRFs by designing decoupled XOR-type function, e.g., $\mathsf{XopEM}$, $\mathsf{SoEM22}$, $\mathsf{SUMPIP}$, and $\mathsf{DS\text{-}SoEM}$ (See Table 1 for a comparison with previous and new quantum attacks on $\mathsf{SoEM22}$). Compared to the previous attack, this offline attack is reduced the query complexity from $\tilde{O}(2^{(n+t)/2})$ to $O(2^{t})$ in the Q2 model. And in the Q1 model, by retaining a tunable truncation parameter $t$, our work significantly optimizes the quantum-time/classical-query tradeoff from $\tilde{O}(2^{2n/3})$, as given by Sun et al. [31], to $\tilde{O}(2^{(2n-t)/3})$, where $t$ is a positive integer.
Organization. The rest of this paper is organized as follows: Section 2 provides the preliminaries, Section 3 discusses the offline dedicated quantum attacks on decoupled XOR-type function, and Section 4 provides a concise conclusion and discussion.
Table 1: Previous and new quantum attacks on $\mathsf{SoEM22}$
Algorithm Model Iteration Oracle query Time Query: c´ Reference
Grover-meets-Simon
algorithm [18] Q2 $2^{n/2}$ $O(2^{n/2}\cdot n)$ $O(2^{n/2}\cdot(n^{3}+T_{f}))$ $2n+\tau+1$ [24]
Dedicated attack
on XOR-type function [23] Q2 $2^{(n-t)/2}$ $O(2^{(n+t)/2}\cdot(n-t))$ $O(2^{(n-t)/2}\cdot(n^{3}+T_{f}))$ $2n-2t+\tau+1$ Ours
Offline dedicated attack
on $p$-XOR-type function Q2 $2^{(n-t)/2}$ $O(2^{t})$ $O(2^{(n-t)/2}\cdot(n^{3}+T_{p}))$ $2n-2t+\tau+1$ Ours
Offline Simon’s algorithm [4] Q1 $2^{2n/3}$ $O(2^{2n/3})$ $O(2^{2n/3}\cdot(n^{3}+T_{f}))$ $2n+\tau+1$ [31]
Offline dedicated attack
on $p$-XOR-type function Q1 $2^{(2n-t)/3}$ $O(2^{(2n-t)/3})$ $O(2^{(2n-t)/3}\cdot(n^{3}+T_{p}))$ $2n-2t+\tau+1$ Ours
  • 1
    The fourth column “Oracle query” indicates classical queries to the encryption oracle in the Q1 model and superposition queries in the Q2 model.
  • 2
  • 3
    $T_{f}$ ($T_{p}$) is the time required to evaluate the function $f$ (the public function $p$) once.
2 Preliminaries
In this section, we review foundational algorithms, including Grover’s algorithm [11], Simon’s algorithm [26], the offline Simon’s algorithm [4], and the dedicated quantum attack on XOR-type function [23].
2.1 Grover’s algorithm
Grover’s algorithm [11] addresses the unstructured search problem, which involves finding a specific element that satisfies a given condition from a database of size $N=2^{n}$. Quantum Amplitude Amplification, as introduced in [5], represents a generalization of Grover’s algorithm.
Problem 1 (Grover’s problem).
@@ 2.1 Grover’s algorithm
2.2 Simon’s algorithm
Simon’s problem involves finding the hidden period $s$ of a two-to-one periodic function. Simon’s algorithm [26] is designed to solve this problem and has an exponential speed advantage over classical algorithms.
Problem 2 (Simon’s Problem).
Given oracle access to $f:\{0,1\}^{n}\to\{0,1\}^{m}$, we promise that there exists a non-trivial $s\neq 0$ such that $f(x)=f(x\oplus s)$ for any $x$, find $s$.
Simon’s algorithm [26] works as follows: First, prepare a uniform superposition $|\phi_{1}\rangle=\frac{1}{\sqrt{2^{n}}}\sum_{x\in\{0,1\}^{n}}|x\rangle|0\rangle$. Then, query the quantum oracle to obtain $f(x)$ in superposition $|\phi_{2}\rangle=\frac{1}{\sqrt{2^{n}}}\sum_{x\in\{0,1\}^{n}}|x\rangle|f(x)\rangle$. And measure the second register. Let $f(z)$ be the measurement result. By omitting the second register, the resulting state is given by $|\phi_{3}\rangle=\frac{1}{\sqrt{2}}(|z\rangle\oplus|z\oplus s\rangle)$. Subsequently, apply Hadamard transformation to $|\phi_{3}\rangle$: $\frac{1}{\sqrt{2^{n+1}}}\sum_{y\in\{0,1\}^{n}}(-1)^{y\cdot z}(1+(-1)^{y\cdot s})|y\rangle$. Finally, measure the first register to obtain a random vector that is orthogonal to $s$. Repeating this process yields a system of linear equations, solved classically to extract $s$. Its key advantage is an exponential quantum speedup with only $O(n)$ queries, making it powerful for cryptanalysis of periodic functions in symmetric cryptography.
2.3 Grover-meets-Simon Algorithm
Grover-meets-Simon algorithm [18] represents a significant breakthrough in quantum cryptanalysis. It ingeniously combines Grover’s algorithm with Simon’s algorithm to solve key-recovery problems in symmetric ciphers.
Problem 3 (Grover-meets-Simon Problem).
Given oracle access to $f:\{0,1\}^{\kappa}\times\{0,1\}^{n}\to\{0,1\}^{m}$, there exists a unique $k_{0}$ such that $f(k_{0},\cdot)$ hides a non-trivial period $s$, the goal is to find $k_{0}$ and the corresponding period $s$.
The algorithm’s framework is based on Grover’s search. It locates the correct key $k_{0}$ by performing a quantum search over the key space $\{0,1\}^{\kappa}$. In each Grover iteration, the algorithm utilizes Simon’s algorithm as a testing function to check whether $f(k,\cdot)$ is periodic or not. Given the output vectors $v_{1},\ldots,v_{c}$ from $c$ parallel Simon routines, it can be verified whether the rank of $\{v_{1},\ldots,v_{c}\}$ is at most n1n-1. Grover-meets-Simon algorithm, which requires $\mathcal{O}(n\cdot 2^{\kappa/2})$ quantum queries and $\mathcal{O}(n^{3}\cdot 2^{\kappa/2})$ time complexity, can be employed against various symmetric cryptographic schemes, such as the FX construction and some MACs.
2.4 The Offline Simon’s Algorithm
In 2019, Bonnetain et al. [4] presented the offline Simon’ algorithm for an asymmetric search of a period problem, with many cryptographic applications.
Problem 4 (Asymmetric Search of a Period Problem).
@@ 2.4 The Offline Simon’s Algorithm
The idea of Algorithm 1 is to decompose the process into two stages: the online phase prepares the superposition state $|\psi_{g}\rangle$ by querying the encryption oracle, and the offline quantum computation phase utilizes $|\psi_{g}\rangle$ to find the target solution via Grover iteration. Algorithm 2 illustrates how to implement the offline procedure, which checks whether the conditional period function $f\oplus g$ has a period without any query to $g$. The online phase is distinguished by the use of superposition queries to the encryption oracle in the Q2 model, whereas only classical queries are available in the Q1 model. The offline quantum computation phase runs in the same way.
Algorithm 1 The Offline Simon’s Algorithm  [4]
@@
Algorithm 2 The procedure test that checks whether $f\oplus g$ has a period, without making any new queries to $g$ [4]
@@
Theorem 1 (Adapted from [4]).
Suppose that $m$ is in $O(n)$, Let $c$ be a sufficiently large constant. Consider Problem 4, there exists the index $i_{0}$ such that $f_{i_{0}}\oplus g$ has a period. Algorithm 1 identifies $i_{0}$ with probability $\Theta(1)$. In the Q2 model, it requires only $O(n)$ quantum queries to $g$ and $O(n2^{m/2})$ quantum queries to $F$. In the Q1 model, it requires $O(2^{n})$ classical queries to $g$ and $O(n2^{m/2})$ quantum queries to $F$. In both cases, the offline computation of Algorithm 1 is done in time $O((n^{3}+nT_{F})2^{m/2})$, where $T_{F}$ is the time needed to evaluate $F$ once.
We now review the specific estimates proposed by Bonnetain et al. [3] for the offline Simon’s algorithm.
Theorem 2 (Adapted from [3]).
Let $f:\{0,1\}^{\kappa}\times\{0,1\}^{n}\to\{0,1\}^{m}$ satisfying Problem 3, $\tau$ be a positive integer and assume that $m\geq\log(4e(n+\kappa+\tau+1))$ and $\kappa\geq 7$, there exists a Grover-meets-Simon attack that uses $\pi/4\text{arcsin}\sqrt{2^{-\kappa}}$ iterations with $c$ quantum queries per iteration. The probability of the success rate being less than
@@
2.5 Dedicated Quantum Attack on XOR-Type Function
XOR-type function constitutes a class of hidden periodic functions with a specific algebraic structure. Specifically, there exists a unique key $k_{0}$ such that the function $f(k_{0},\cdot)$ possesses a hidden period $s$, satisfying: $f(k_{0},x)=f(k_{0},x\oplus s)$. Furthermore, the function can be decomposed into two components: one independent of the key $k$, denoted as $g_{1}(x)$, and the other dependent on both the input $x$ and the key $k$, denoted as $g_{2}(k,x)$. This leads to the following expression for XOR-type function:

### v3

# Offline Dedicated Quantum Attacks on Block Cipher Constructions Based on Two Parallel Permutation-Based Pseudorandom Functions

@@ Abstract
Quantum cryptanalysis is essential for evaluating the security of cryptographic systems against the threat of quantum computing. Recently, Shi et al. introduced a dedicated quantum attack on block cipher constructions based on XOR-type functions, which greatly reduces the required resources (including circuit depth, width, and the number of gates) compared to the parallel Grover-meets-Simon algorithm. Here, our contribution is in two aspects. On the one hand, we discover new cryptographic structures amenable to this attack: PolyMAC and constructions based on two parallel permutation-based pseudorandom functions (TPP-PRFs), including $\mathsf{XopEM}$, $\mathsf{SoEM22}$, $\mathsf{SUMPIP}$, and $\mathsf{DS\text{-}SoEM}$, thereby answering Shi et al.’s open question. On the other hand, for constructions based on TPP-PRFs, we break the obstacle that this attack relies on online query by constructing decoupled XOR-type functions, then propose an offline quantum attack on them. Compared to previous results, our offline attack exhibits significantly reduced query complexity. Specifically, the number of queries to the encryption oracle is reduced from $O(2^{(n+t)/2}\cdot(n-t))$ to $O(2^{t}\cdot(n-t))$ in the quantum query model, where $0<t<n$, $t$ is a truncation parameter, and $n$ is the input length of constructions. Further, we enable its implementation in the classical query model, optimizing both the classical query complexity and time complexity from $\tilde{O}(2^{2n/3})$ to $\tilde{O}(2^{(2n-t)/3})$.
publicationid: pubid:
I Introduction
Quantum computing, rooted in the principles of quantum mechanics, exhibits speed-up advantages for specific computational problems, such as linear system solving [harrow2009quantum, wan2018asymptotic], quantum simulation [Li2025quantumsimulation], and machine learning [li2025efficient, huang2021power]. As a special kind of quantum algorithm for Noisy Intermediate Scale Quantum devices, the Variational Quantum Algorithm (VQA) [cerezo2021variational, wang2021noise, su2025topology] is currently being extensively studied, showing potential advantages in tasks such as combinatorial optimization [wu2025resource, wu2025], image classification [wu2024quantum, sein2025image], and so on [peruzzo2014variational, song2024quantum].
The development of quantum computing poses unprecedented threats to cryptographic security. Notably, Shor’s algorithm [shor1994algorithms] and Grover’s algorithm [grover1996fast] fundamentally threaten public-key and symmetric cryptography, respectively. Meanwhile, VQAs have also garnered attention in quantum cryptanalysis [wang2025reducing, wang2022variational], particularly for analyzing hard problems underlying post-quantum cryptography, such as Shortest Vector Problem [albrecht2023variational], Learning With Errors Problem [zheng2025quantum, zeng2025analysis].
@@ I Introduction
Simon’s algorithm [simon1997power] provides a crucial tool for quantum cryptanalysis by achieving exponential speedup over classical methods in finding the period of two-to-one periodic functions. This capability directly threatens the security of block ciphers (e.g., $3$-round Feistel [kuwakado2010quantum], Even-Mansour cipher [kuwakado2012security, shinagawa2022quantum, zhang2023quantum], etc.). In 2017, Grover-meets-Simon algorithm, which is presented by Leander and May [leander2017grover], demonstrated that key whitening does not enhance security in the quantum-chosen-plaintext attack (q-CPA) setting on the FX construction. Subsequently, Grover-meets-Simon algorithm has played a pivotal role in researching the quantum security of symmetric cryptography, including $\mathsf{Feistel}$ [dong2018quantum, ito2019quantum], $\mathsf{Lai\text{-}Massey}$ [mao2022quantum], $\mathsf{MISTY}$ structure [canale2022simon], $\mathsf{SUM\text{-}ECBC}$ [guo2021attacks], and several Beyond-Birthday-Bound MACs [sun2023quantum].
@@
The above attacks, based on Simon’s and Grover-meets-Simon algorithms, rely on superposition queries to the encryption oracle (the Q2 model), presenting significant implementation challenges. This has spurred growing interest in a quantum attack that utilizes classical queries and offline quantum computation (the Q1 model) in recent years. In 2019, Bonnetain et al. [bonnetain2019quantum] proposed the offline Simon’s algorithm and achieved key-recovery attacks on the FX and Even-Mansour constructions in the Q1 model. Following this, Li et al. [li2025quantum] proposed a quantum key recovery attack on $\mathsf{SoEM21}$ and $\mathsf{SoEMs1}$ in the Q1 model. Further, Sun et al. [sun2025quantum] systematically studied the quantum security of block cipher constructions based on permutation-based pseudorandom functions in the Q1 model.
@@
Due to the limitations of quantum computer implementation technology, reducing the circuit depth of quantum attacks is essential for the practical implementation. Shi et al. [shi2024dedicated] proposed a dedicated quantum attack on constructions based on XOR-type functions of the form $f(\alpha_{k},x)=g_{1}(x)\oplus g_{2}(\alpha_{k}\oplus x)$. In contrast to previous attacks, their attack reduced circuit costs, including depth, width, and gate count. Moreover, they applied this attack to several Beyond-Birthday-Bound (BBB) MACs, such as $\mathsf{SUM\text{-}ECBC}$, its variants $2\mathsf{K}$-$\mathsf{SUM\text{-}ECBC}$, $2\mathsf{K}$-$\mathsf{ECBC\_Plus}$, and $\mathsf{GCM\text{-}SIV2}$. For $\mathsf{SUM\text{-}ECBC}$, it simultaneously achieved savings of at least $62.3\%$ in depth, $19.5\%$ in width, and $22.2\%$ in gate count.
@@
Motivations. Our work is motivated by two critical questions arising from the above dedicated attack.
@@
First, the dedicated quantum attack proposed by Shi et al. [shi2024dedicated] is currently limited to the BBB MACs mentioned above. They posed an open question that is to identify other cryptographic constructions that are vulnerable to this attack.
@@
Second, this attack on these BBB MACs relies on quantum queries to the encryption oracle. This naturally leads to a more challenging question, that is, can this quantum attack be performed in the offline setting?
@@
Our Contributions. Addressing these challenges of applicability and practicality, our work achieves the following results:
@@
    We construct XOR-type functions based on PolyMAC and two parallel permutation-based pseudorandom functions (TPP-PRFs), that is, we identify new instances that are vulnerable to the attack presented by Shi et al. [shi2024dedicated], thereby answering their open question.
@@
    We propose an OFFLINE dedicated quantum attack on constructions based on TPP-PRFs by designing decoupled XOR-type functions, successfully applying it to constructions like $\mathsf{XopEM}$, $\mathsf{SoEM22}$, $\mathsf{SUMPIP}$, and $\mathsf{DS\text{-}SoEM}$. Compared to the previous attack, the query complexity in our attack is reduced from $O(2^{(n+t)/2}\cdot(n-t))$ to $O(2^{t}\cdot(n-t))$ in the Q2 model, where $0<t<n$, $t$ is a truncation parameter, and $n$ is the input length of constructions. Furthermore, we significantly optimize the time complexity and classical query complexity from $\tilde{O}(2^{2n/3})$, as given by Sun et al. [sun2025quantum], to $\tilde{O}(2^{(2n-t)/3})$ in the Q1 model.
Organization. The rest of this paper is organized as follows: Section II provides the preliminaries, Section III discusses an offline dedicated quantum attack on block ciphers by constructing decoupled XOR-type functions, and Section IV provides a concise conclusion and discussion.
II Preliminaries
In this section, we review foundational algorithms, including Grover’s algorithm [grover1996fast], Simon’s algorithm [simon1997power], the offline Simon’s algorithm [bonnetain2019quantum], and the dedicated quantum attack on constructions based on XOR-type functions [shi2024dedicated].
II-A Grover’s algorithm
Grover’s algorithm [grover1996fast] addresses the unstructured search problem, which involves finding a specific element that satisfies a given condition from a database of size $N=2^{n}$. Quantum Amplitude Amplification, introduced in [brassard2000quantum], is a generalization of Grover’s algorithm.
Problem 1 (Grover’s problem)
@@ II-A Grover’s algorithm
II-B Simon’s algorithm
Simon’s problem involves finding the unique hidden period $s$ of a two-to-one periodic function. Simon’s algorithm [simon1997power] solves this problem and provides an exponential speed advantage over classical algorithms.
Problem 2 (Simon’s Problem)
Given oracle access to a function $f:\{0,1\}^{n}\to\{0,1\}^{m}$, we are promised that there exists a non-trivial period $s\neq 0$ such that $f(x)=f(x\oplus s)$ for any $x\in\{0,1\}^{n}$. The goal is to find $s$.
Simon’s algorithm [simon1997power] works as follows: First, prepare a uniform superposition state $|\phi_{1}\rangle=\frac{1}{\sqrt{2^{n}}}\sum_{x\in\{0,1\}^{n}}|x\rangle|0\rangle$. Then, query the quantum oracle to obtain $f(x)$ in superposition $|\phi_{2}\rangle=\frac{1}{\sqrt{2^{n}}}\sum_{x\in\{0,1\}^{n}}|x\rangle|f(x)\rangle$. Then, measure the second register. Let the measurement outcome be $f(z)$. By tracing out the second register, the first register collapses to a superposition $|\phi_{3}\rangle=\frac{1}{\sqrt{2}}(|z\rangle+|z\oplus s\rangle)$. Subsequently, apply the Hadamard transformation to $|\phi_{3}\rangle$: $\frac{1}{\sqrt{2^{n+1}}}\sum_{y\in\{0,1\}^{n}}(-1)^{y\cdot z}(1+(-1)^{y\cdot s})|y\rangle$. Finally, measure the first register to obtain a random vector $y$ such that $y\cdot s=0\mod 2$. Repeating this process yields a system of linear equations, which is then solved classically to efficiently extract $s$. The key advantage of Simon’s algorithm is an exponential quantum speedup with only $O(n)$ queries, making it a powerful tool for the cryptanalysis of periodic functions in symmetric cryptography.
II-C The Offline Simon’s Algorithm
In 2019, Bonnetain et al. [bonnetain2019quantum] presented the offline Simon’s algorithm to solve the asymmetric search of a period problem, which has some cryptographic applications.
Problem 3 (Asymmetric Search of a Period Problem)
@@ II-C The Offline Simon’s Algorithm
Algorithm 1 The Offline Simon’s Algorithm  [bonnetain2019quantum]
@@
Algorithm 2 The procedure test that checks whether $f\oplus g$ has a period, without making any new queries to $g$ [bonnetain2019quantum]
@@
The idea of Algorithm 1 is to decompose the process into two stages: the online phase prepares the superposition state $|\psi_{g}\rangle$ by querying the encryption oracle, and the offline quantum computation phase utilizes $|\psi_{g}\rangle$ to find the target solution using the Simon procedure. Algorithm 2 illustrates how to implement the offline procedure, which checks whether the conditional period function $f\oplus g$ has a period without any query to $g$. The Q1 and Q2 models fundamentally differ only in the online phase. The superposition queries to the encryption oracle are permitted in the Q2 model, whereas the classical queries are restricted in the Q1 model. The subsequent offline quantum computation phase remains identical in both models.
@@
Theorem 1
(Adapted from [bonnetain2019quantum]) Assume that $m=O(n)$, and let $c$ be a sufficiently large constant. Consider Problem 3, there exists a unique index $i_{0}$ such that $f_{i_{0}}\oplus g$ has a period. Algorithm 1 identifies $i_{0}$ with probability $\Theta(1)$. In the Q2 model, it requires only $O(n)$ quantum queries to $g$ and $O(n2^{m/2})$ quantum queries to $F$. In the Q1 model, it requires $O(2^{n})$ classical queries to $g$ and $O(n2^{m/2})$ quantum queries to $F$. In both cases, the offline computation time of Algorithm 1 is $O((n^{3}+nT_{F})2^{m/2})$, where $T_{F}$ is the time needed to evaluate $F$ once.
We now review the specific estimates proposed by Bonnetain [bonnetain2021tight] for the offline Simon’s algorithm.
Theorem 2
(Adapted from [bonnetain2021tight]) Let $f:\{0,1\}^{\kappa}\times\{0,1\}^{n}\to\{0,1\}^{m}$, and assume that there exists a unique $k_{0}$ such that $f(k_{0},x)=f(k_{0},x\oplus s)$, $\tau$ be a positive integer, $m\geq\log(4e(n+\kappa+\tau+1))$, and $\kappa\geq 7$. There exists the offline Simon attack that uses $\pi/4\text{arcsin}\sqrt{2^{-\kappa}}$ iterations with $c$ quantum queries per iteration. The probability of the success rate being less than
@@
II-D Dedicated Quantum Attack on Constructions Based on XOR-Type Function
The XOR-type function is a class of hidden periodic functions with a specific algebraic structure. For a function $f$, there exists a unique key $k_{0}$ such that $f(k_{0},x)=f(k_{0},x\oplus s)$. Moreover, $f(k,x)$ can be decomposed into two components: $g_{1}(x)$, independent of the key $k$, and $g_{2}(x\oplus\alpha_{k})$, dependent on the input $x$ and the secret state $\alpha_{k}$. The XOR-type function is expressed as